OpenStudy (anonymous):

what fraction falls between 5/14 and 5/7

OpenStudy (anonymous):

infinitely many, lol. but to give you some multiples of 5: \[\frac{ 5 }{ 13 },\frac{ 5 }{ 12 }, \frac{ 5 }{ 11 }, \frac{ 5 }{ 10 }, \frac{ 5 }{ 9 }, \frac{ 5 }{ 8 }\]